Circular No. 6955

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cq IN ANONYMOUS GALAXY L. Germany reports that the Mount Stromlo Abell Cluster Supernova

Search Team (cf. IAUC 6639) has discovered another apparent supernova (V = 19.3) on V and R CCD images taken on June 17 (limiting mag V = 20.5) by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera), and confirmed by Chan on June 18 (limiting mag V = 20.5). SN 1998cq is located near Abell 2819 at α = 0^{h}45^{m}06^{s}.11, δ = −63^{o}49′50′′.4 (equinox 2000.0), which is 0′′.3 east and 1′′.2 north of the center of its host galaxy. A nearby star has position end figures 03^{s}.14, 14′′.8. SN 1998cq was not visible on similar images taken on 1997 Dec. 19 (limiting mag V = 21.5). A spectrum (range 400–750 nm) taken by Germany with the Mount Stromlo 2.3-m telescope (+ DBS) on June 26 is consistent with a type-Ia supernova at a redshift of 0.15. Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

XTE J2123–058 A. Levine, Massachusetts Institute of Technology (MIT); and J. Swank

and E. Smith, Goddard Space Flight Center (GSFC), NASA, report on behalf of the RXTE ASM and PCA teams at MIT and GSFC: “A transient x-ray source has been detected with the RXTE/ASM at α = 21^{h}23^{m}.1, δ = −5^{o}49′ (equinox 2000.0), corresponding to l = 46^{o}.4, b = −36^{o}.2. The position uncertainty is ∼ 10′ along an axis with p.a. 20^{o} (i.e., northeastsouthwest) and ∼ 5′ in the perpendicular direction (90-percent confidence). The flux was measured to be ∼ 100 mCrab (2–12 keV) in each of five measurements on June 27–28. The strength of the source has been confirmed with PCA observations June 27.95, in which the count rate varied from 600 to 1000 counts/s. Additional PCA observations, including a determination of the source location, will be undertaken. The appearance of a new transient x-ray source at high galactic latitude is unusual; we urge observers to carry out additional observations at all wavelengths.”

NOVA OPHIUCHI 1998 Additional CCD photometry by D. Hanžl, Brno, Czech Republic (cf.

IAUC 6943): June 19.887 UT, V = 10.87 ± 0.02, B − V = +0.75 ± 0.03; 20.931, 10.86 ± 0.02, +0.67 ± 0.03; 21.928, 11.87 ± 0.01, +0.54 ± 0.02; 22.878, 12.22 ± 0.06, +0.54 ± 0.02; 25.942, 12.92 ± 0.04, +0.64 ± 0.1; 27.891, 13.39 ± 0.07, –.

1998 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green